"In November 2005, the exchange rate of Canadian dollars to U.S dollars was 1 to 0.85. A Canadian citizen paid $12.28 in U.S dollars for lunch while visiting New York City. If the exchange rate were 1.28 Canadian dollars to 1 U.S dollar, what would the cost have been in Canadian dollars?"
Answer by TimothyLamb(4379)

ca/us = 1.28/1
---
ca/us = 1.28/1
ca/12.28 = 1.28
ca = 1.28*12.28 = CA$15.72
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php